# German General Personas: A Survey-Derived Persona Prompt Collection for Population-Aligned LLM Studies

### v1

@@ Abstract
The use of Large Language Models (LLMs) for simulating human perspectives via persona prompting is gaining traction in computational social science. However, well-curated, empirically grounded persona collections remain scarce, limiting the accuracy and representativeness of such simulations. Here we introduce the German General Personas (GGP) collection, a comprehensive and representative persona prompt collection built from the German General Social Survey (ALLBUS). The GGP and its persona prompts are designed to be easily plugged into prompts for all types of LLMs and tasks, steering models to generate responses aligned with the underlying German population. We evaluate GGP by prompting various LLMs to simulate survey response distributions across diverse topics, demonstrating that GGP-guided LLMs outperform state-of-the-art classifiers, particularly under data scarcity. Furthermore, we analyze how the representativity and attribute selection within persona prompts affect alignment with population responses. Our findings suggest that GGP provides a potentially valuable resource for research on LLM-based social simulations that enables more systematic explorations of population-aligned persona prompting in NLP and social science research.
@@
German General Personas: A Survey-Derived Persona Prompt Collection for Population-Aligned LLM Studies
Jens Rupprecht1, Leon Fröhling2, Claudia Wagner2,3,4, Markus Strohmaier1,2,4
@@ 1.  Introduction
Simulating human behavior represents a complex challenge given the vast diversity of human experiences and perspectives. Traditional research methods, such as large-scale question-oriented surveys, are crucial for informing social science research and policy decisions, but they are often costly and time-consuming to administer. As a result, researchers are increasingly turning to innovative approaches to capture the nuances of human behavior. One recent avenue is the use of large language models (LLMs) and so-called persona prompting chen2024from; tseng2024two; lutz2025prompt; zhang2025personalization, i.e., the use of personas descriptions to induce role-playing in LLMs, as an accessible option to steer and control simulated human behavior in surveys ma2025algorithmic; miranda2025simulating.
@@
Persona prompting refers to a broad range of prompting techniques that aim to harness the general world knowledge captured by LLMs and steer the generation of simulated survey responses toward the perspective of the persona introduced in the prompt. Compared to alternative approaches for the alignment of LLMs, such as training dedicated models feng2023pretraining or fine-tuning existing ones orlikowski2025beyond; suh2025language, prompt-based approaches offer a number of crucial advantages: they are modular (i.e., they can be used with different models and for different tasks), upgradable (i.e., they can be modified and can be expected to improve in performance with the release of improved models), and shareable (i.e., the prompts can easily be made available to other potential uses and users), among other advantages.
@@
Figure 1: Grounding the German General Personas (GGP) in the ALLBUS survey. We construct individual persona prompts for each ALLBUS participant, varying size and composition via available attributes. A global variable importance ranking informs the selection of the $k$ most important attributes (TOP-$k$). Personas comprise a fixed block of core socio-demographics and a more extensible block of TOP-$k$ attributes that allow varying information content. The GGP is available in JSON- and full-text formats. The originally German survey items and personas are here presented in English for illustration.
@@
German General Personas are a prompt collection that consists of $5,246$ personas constructed from a representative sample of the German population. By grounding the German General Personas collection in the information available in the ALLBUS (allbus2023)—the German General Social Survey and thus the premier source of information on the attributes and attitudes of the German population—we demonstrate that more systematic and principled approaches to the construction of persona prompt collections are possible.
@@
GGP offers various potential advantages: First, personas introduce relevant contextual information to the model, providing it with the necessary information to meaningfully anchor predictions for a task or target variable in empirically observed associations and connections. Second, the ALLBUS is a probability-based survey and the personas derived from it aim to represent the German population. While there has been a focus on the biased representation of populations that LLMs default towards in their generation of survey responses santurkar2023whose; durmus2024towards; sen2025missing, the GGP may potentially help to align LLMs better with the German population.
@@
As a main contribution, this paper introduces German General Personas  [^1] German General Personas Github Repository: https://github.com/germanpersonas/German-General-Personas.git—a novel textual resource for NLP and Computational Social Science (CSS). We describe the creation process of GGP and offer a preliminary analysis of its potential and utility through two exemplary demonstrations.
@@ 2.  Related Work
Persona prompting is a versatile and accessible method for social simulations and pluralistic alignment with LLMs anthis2025position; sorensen2024position. There is already a wide range of applications and promising demonstrations across disciplines and tasks, including (but not limited to): vote choice prediction argyle2023out; von2025vox, generation of synthetic public opinions hwang2023aligning; ma2025algorithmic; miranda2025simulating, annotation of (subjective) constructs hu2024quantifying; beck2024sensitivity, simulations of participants in economic and social science experiments aher2023using; hewitt2024predicting, including for underrepresented and otherwise hard to reach populations gonzalez2025llms, simulation of pluralistic debates ashkinaze2025plurals, behavioral simulation of users chen2025personatwin; salem2025tinytroupe, and even productive real world tasks such as red teaming deng2025personateaming. Systematic reviews of the literature on the use of LLMs for role-playing and persona prompting provide an overview of this dynamic area of research lutz2025prompt; zhang2025personalization; chen2024from; tseng2024two.
@@
Beyond purely prompt-based approaches, kim2023ai propose a framework that fine-tunes LLMs on survey data for predicting responses at an individual level, and cao2025specializing and suh2025language use group-level information to fine-tune LLMs for the simulation of response distributions. orlikowski2025beyond compare the use of personas for (zero-shot) persona prompting with fine-tuning LLMs on pairs of annotators’ persona descriptions and their annotations, finding that fine-tuning strictly outperforms zero-shot in simulating individual-level annotations. Apart from fine-tuning, chen2025persona introduce the idea of persona vectors, extracted via persona descriptions and used to monitor and steer LLM generations. sun2025persona propose an approach that leverages personas via retrieval-augmentation to personalize LLM responses.
@@
Conceptual criticism of the use of persona prompting include the work by cheng2023marked; cheng2023compost, who show that simulations of political and marginalized groups tend towards caricature, and that LLM-generated personas of non-white, non-male demographics exhibit patterns of othering and exoticization. Similarly, wang2025large show how LLMs are likely to misportray and flatten the representation of demographic groups. kim2024persona investigate how the use of personas might degrade performance by distracting the model from the task at hand. kirk2024benefits discuss the broader context and societal implications of aligning LLMs with individuals.
@@
In the literature, there are some noteworthy demonstrations of how well-curated, empirically-grounded persona collections lead to improved performance in their applications. park2024generative show how the use of transcripts of hour-long interviews with participants, i.e., extremely information-rich persona context, leads to impressive LLM performance in predicting individuals’ survey responses. Similarly, toubia2025database survey a representative sample of US-based crowdworkers to collect extensive information on their demographic-, psychological-, economic-, personality-, and cognitive measures, and peng2025mega show that the use of this detailed personal information improves the correlations between human and LLM-simulated responses.
@@
However, this quality comes at the price of up to \$$100$ paid to $1,052$ participants for hour-long interviews and \$$37$ paid to $2,058$ crowdworkers for answering $500$ survey questions, respectively. Using freely and publicly available survey data offers an accessible alternative for grounding persona collections in high-quality empirical information. Most similar to our work, castricato2025persona introduce PERSONA, a collection of $1,586$ synthetic personas created from US census data. In contrast to their approach, our GGP establishes 1:1 correspondence between survey respondents and persona prompts, and automatically identifies important persona attributes from the full range of survey variables instead of limiting selection to a set of $31$ variables.
@@
With the introduction of GGP, we aim to contribute a well-curated, empirically-grounded persona collection as a crucial resource for meaningful advances in research on persona-based LLM personalization.
3.  German General Personas
@@
German General Personas is derived from the ALLBUS, the German General Social Survey (Figure 1). While the GGP is necessarily tied to the German population at a specific moment in time, we propose a creation process that can be applied to different (population) surveys in order to create survey-derived persona collections for different contexts and purposes.
Grounding German General Personas
@@
Ideally, a population survey panel used for the construction of persona prompt collections should cover a set of core demographics and additional attributes for a representative sample drawn from the population of interest. Following the discussion on the use of the terms representative sample and representativity in the scientific literature by kruskal1979representative, we believe that population-level general social surveys such as the ALLBUS in Germany, the General Social Survey (GSS) in the US, the British Social Attitudes Survey (BSA) in the UK, or the European Social Survey (ESS) covering 31 European Countries in its latest round, could represent plausible empirical foundations for representative persona prompt collections.
@@
To create this first version  [^2] We clearly identify our persona collection by referencing the source and year: GGP-ALLBUS-2023. of the GGP, we use the latest release of the biannual ALLBUS (allbus2023), released in January 2025 and including responses collected between April and September 2023. It features $5,246$ participants and a total of $605$ variables. We work with the ALLBUScompact  [^3] https://search.gesis.org/research_data/ZA8832 (allbus2023compact), the openly available version of the ALLBUS, featuring $579$ variables. For data protection reasons, the ALLBUScompact contains information in reduced detail on the socio-demographic characteristics age, origin, occupation and income, and omits all fine-grained geographic variables.
@@
In our preliminary analysis, we explore the effects of using a representative persona collection such as the GGP—representative of the German population by virtue of mirroring all ALLBUS participants—as well as unrepresentative variants of it, created by systematically oversampling on selected attributes (Section 4).
@@
While it is theoretically feasible to use all $579$ variables featured in the latest release of the ALLBUScompact in the persona prompts, this would unavoidably come at the price of increased context window requirements, computational resources, as well as potentially degrading performance due to the dilution of relevant information or the failure of the model to put equal attention to all variables in a lengthy persona description. For example, shi2023large have shown that LLMs can easily be distracted by irrelevant information included in the context window, and de2025principled have observed performance drops after including irrelevant persona details in persona prompts.
One of the core advantages of general population surveys such as the ALLBUS is the coverage of a variety of different topics, ranging from core demographic variables such as age and gender to very specific, thematic variables covering, e.g., the participant’s political views or lifestyle choices. Therefore, we design a data-driven method that allows us to identify and select the statistically most important variables, instead of having to (artificially) restrict the GGP to cover only task-specific areas such as socio-demographics, psychological inventories, or political beliefs.
@@
We expect that variables with higher aggregate feature importance score explain more variance of the other variables in the survey. This ranking thus allows us to create versions of the GGP collection for different computational budgets and needs by selecting only the $k$ most important attributes (TOP-$k$). In situations with restricted prompt length or inference time, versions with less attributes may be preferable, while in unrestricted settings, the number of included attributes may be increased, up to the inclusion of all $380$ survey variables. In our preliminary analysis, we explore the effects of increasing the number of persona attributes on the ability of LLMs to accurately simulate survey response distributions (Section 4).
An important limitation of our approach is its inability to deal with missing values. Because large surveys such as the ALLBUS work with questionnaire splits to manage the number of questions each participant is asked to answer, no participant will have answered all survey questions. Thus, some information is just not available for some of the participants, meaning that for any given variable some persona descriptions will always be incomplete, missing individual attributes. Table 1 shows the extent to which the persona descriptions across the different TOP-$k$ collections are complete. The $k$ thus represents the maximum number of attributes possibly featured in a persona of a given TOP-$k$ collection, and the reported average is the actual number of attributes included. We discuss the implications for the GGP and our preliminary analysis in Section 6.
@@
There is a broad range of different prompting techniques in the literature lutz2025prompt, ranging from the use of key-value pairs in a dictionary or JSON-format (e.g., castricato2025persona) to the use of full-text persona descriptions (e.g., ge2024scaling), oftentimes generated using LLMs. While personas in the JSON-format are easier to construct and more convenient to work with, the full-text persona descriptions resemble the natural language interactions that especially instruction-tuned LLMs have been optimized for more closely.
We evaluate the impact of using different persona formats for predicting survey response distributions. Since we do not observe a large impact of the persona format on the experimental results, (see Appendix A Figure 7) we use the JSON-format for all following experiments, expecting the results to generalize to other persona formats. However, we make both the JSON as well as the full-text version of the GGP publicly available. We validate the full-text personas that are LLM-generated from the JSON-personas as input, manually checking a sample of $80$ personas ($20$ randomly-sampled personas per TOP-$2$, -44, -88 and -$16$ collection) for hallucinations, misrepresentation or omission of individual attributes. We find that $69$ of $80$ (or 86.2586.25%) full-text personas are accurate depictions of the JSON-personas, with errors occurring rarely through misrepresentation, e.g., of double negations or complex temporal orders, or through the omission of single attributes.

### v2

# arXiv:2511.21722v2[cs.CL] 02 Mar 2026

@@ Abstract
The use of Large Language Models (LLMs) for simulating human perspectives via persona prompting is gaining traction in computational social science. However, well-curated, empirically grounded persona collections remain scarce, limiting the accuracy and representativeness of such simulations. Here, we introduce the German General Social Survey Personas (GGSS Personas) collection, a comprehensive and representative persona prompt collection built from the German General Social Survey (ALLBUS). The GGSS Personas and their persona prompts are designed to be easily plugged into prompts for all types of LLMs and tasks, steering models to generate responses aligned with the underlying German population. We evaluate GGSS Personas by prompting various LLMs to simulate survey response distributions across diverse topics, demonstrating that GGSS Personas-guided LLMs outperform state-of-the-art classifiers, particularly under data scarcity. Furthermore, we analyze how the representativity and attribute selection within persona prompts affect alignment with population responses. Our findings suggest that GGSS Personas provide a potentially valuable resource for research on LLM-based social simulations that enables more systematic explorations of population-aligned persona prompting in NLP and social science research.
@@
German General Social Survey Personas: A Survey-Derived Persona Prompt Collection for Population-Aligned LLM Studies
Jens Rupprecht1,*, Leon Fröhling2,*, Claudia Wagner2,3,4, Markus Strohmaier1,2,4
@@ 1.  Introduction
Simulating human behavior represents a complex challenge given the vast diversity of human experiences and perspectives. Traditional research methods, such as large-scale question-oriented surveys, are crucial for informing social science research and policy decisions, but they are often costly and time-consuming to administer. As a result, researchers are increasingly turning to innovative approaches to capture the nuances of human behavior. One recent avenue is the use of large language models (LLMs) and so-called persona prompting Chen et al. (2024); Tseng et al. (2024); Lutz et al. (2025); Zhang et al. (2025), i.e., the use of personas descriptions to induce role-playing in LLMs, as an accessible option to steer and control simulated human behavior in surveys Ma et al. (2025); Miranda and Balbi (2025).
@@
Persona prompting refers to a broad range of prompting techniques that aim to harness the general world knowledge captured by LLMs and steer the generation of simulated survey responses toward the perspective of the persona introduced in the prompt. Compared to alternative approaches for the alignment of LLMs, such as training dedicated models Feng et al. (2023) or fine-tuning existing ones Orlikowski et al. (2025); Suh et al. (2025), prompt-based approaches offer a number of crucial advantages: they are modular (i.e., they can be used with different models and for different tasks), upgradable (i.e., they can be modified and can be expected to improve in performance with the release of improved models), and shareable (i.e., the prompts can easily be made available to other potential uses and users), among other advantages.
@@
Figure 1: Grounding the German General Social Survey Personas (GGSS Personas) in the ALLBUS survey. We construct individual persona prompts for each ALLBUS participant, varying size and composition via available attributes. A global variable importance ranking informs the selection of the $k$ most important attributes (TOP-$k$). Personas comprise a fixed block of core socio-demographics and a more extensible block of TOP-$k$ attributes that allow varying information content. The GGSS Personas is available in JSON- and full-text formats. The originally German survey items and personas are presented in English for illustration.
@@
German General Social Survey Personas are a prompt collection that consists of $5,246$ personas constructed from a representative sample of the German population. By grounding the German General Social Survey Personas collection in the information available in the ALLBUS (GESIS, 2025a)—the German General Social Survey and thus the premier source of information on the attributes and attitudes of the German population—we demonstrate that more systematic and principled approaches to the construction of persona prompt collections are possible  [^1] German General Social Survey Personas are available after registration: https://doi.org/10.4232/1.14707.
@@
GGSS Personas offers various potential advantages: First, personas introduce relevant contextual information to the model, providing it with the necessary information to meaningfully anchor predictions for a task or target variable in empirically observed associations and connections. Second, the ALLBUS is a probability-based survey and the personas derived from it aim to represent the German population. While there has been a focus on the biased representation of populations that LLMs default towards in their generation of survey responses Santurkar et al. (2023); Durmus et al. (2024); Sen et al. (2025), the GGSS Personas may potentially help to align LLMs better with the German population.
@@
As a main contribution, this paper introduces German General Social Survey Personas—a novel textual resource for NLP and Computational Social Science (CSS). We describe the creation process of GGSS Personas and offer a preliminary analysis of its potential and utility through two exemplary demonstrations.
@@ 2.  Related Work
Persona prompting is a versatile and accessible method for social simulations and pluralistic alignment with LLMs Anthis et al. (2025); Sorensen et al. (2024). There is already a wide range of applications and promising demonstrations across disciplines and tasks, including (but not limited to): vote choice prediction Argyle et al. (2023); von der Heyde et al. (2025), generation of synthetic public opinions Hwang et al. (2023); Ma et al. (2025); Miranda and Balbi (2025), annotation of (subjective) constructs Hu and Collier (2024); Beck et al. (2024); Fröhling et al. (2025), simulations of participants in economic and social science experiments Aher et al. (2023); Hewitt et al. (2024), including for underrepresented and otherwise hard to reach populations Gonzalez-Bonorino et al. (2025), simulation of pluralistic debates Ashkinaze et al. (2025), behavioral simulation of users Chen et al. (2025b); Salem et al. (2025), and even productive real world tasks such as red teaming Deng et al. (2025). Systematic reviews of the literature on the use of LLMs for role-playing and persona prompting provide an overview of this dynamic area of research Lutz et al. (2025); Zhang et al. (2025); Chen et al. (2024); Tseng et al. (2024).
@@
Beyond purely prompt-based approaches, Kim and Lee (2023) propose a framework that fine-tunes LLMs on survey data for predicting responses at an individual level, and Cao et al. (2025) and Suh et al. (2025) use group-level information to fine-tune LLMs for the simulation of response distributions. Orlikowski et al. (2025) compare the use of personas for (zero-shot) persona prompting with fine-tuning LLMs on pairs of annotators’ persona descriptions and their annotations, finding that fine-tuning strictly outperforms zero-shot in simulating individual-level annotations. Apart from fine-tuning, Chen et al. (2025a) introduce the idea of persona vectors, extracted via persona descriptions and used to monitor and steer LLM generations. Sun et al. (2025) propose an approach that leverages personas via retrieval-augmentation to personalize LLM responses.
@@
Conceptual criticism of the use of persona prompting include the work by Cheng et al. (2023a, b), who show that simulations of political and marginalized groups tend towards caricature, and that LLM-generated personas of non-white, non-male demographics exhibit patterns of othering and exoticization. Similarly, Wang et al. (2025) show how LLMs are likely to misportray and flatten the representation of demographic groups. Kim et al. (2024) investigate how the use of personas might degrade performance by distracting the model from the task at hand. Kirk et al. (2024) discuss the broader context and societal implications of aligning LLMs with individuals.
@@
In the literature, there are some noteworthy demonstrations of how well-curated, empirically-grounded persona collections lead to improved performance in their applications. Park et al. (2024) show how the use of transcripts of hour-long interviews with participants, i.e., extremely information-rich persona context, leads to impressive LLM performance in predicting individuals’ survey responses. Similarly, Toubia et al. (2025) survey a representative sample of US-based crowdworkers to collect extensive information on their demographic-, psychological-, economic-, personality-, and cognitive measures, and Peng et al. (2025) show that the use of this detailed personal information improves the correlations between human and LLM-simulated responses.
@@
However, this quality comes at the price of up to \$$100$ paid to $1,052$ participants for hour-long interviews and \$$37$ paid to $2,058$ crowdworkers for answering $500$ survey questions, respectively. Using freely and publicly available survey data offers an accessible alternative for grounding persona collections in high-quality empirical information. Most similar to our work, Castricato et al. (2025) introduce PERSONA, a collection of $1,586$ synthetic personas created from US census data. In contrast to their approach, our GGSS Personas establishes 1:1 correspondence between survey respondents and persona prompts, and automatically identifies important persona attributes from the full range of survey variables instead of limiting selection to a set of $31$ variables.
@@
With the introduction of GGSS Personas, we aim to contribute a well-curated, empirically-grounded persona collection as a crucial resource for meaningful advances in research on persona-based LLM personalization.
3.  German General Social Survey Personas
@@
German General Social Survey Personas is derived from the ALLBUS, the German General Social Survey (Figure 1). While the GGSS Personas is necessarily tied to the German population at a specific moment in time, we propose a creation process that can be applied to different (population) surveys in order to create survey-derived persona collections for different contexts and purposes.
Grounding German General Social Survey Personas
@@
Ideally, a population survey panel used for the construction of persona prompt collections should cover a set of core demographics and additional attributes for a representative sample drawn from the population of interest. Following the discussion on the use of the terms representative sample and representativity in the scientific literature by Kruskal and Mosteller (1979), we believe that population-level general social surveys such as the ALLBUS in Germany, the General Social Survey (GSS) in the US, the British Social Attitudes Survey (BSA) in the UK, or the European Social Survey (ESS) covering 31 European Countries in its latest round, could represent plausible empirical foundations for representative persona prompt collections.
@@
To create this first version of the GGSS Personas, we use the latest release of the biannual ALLBUS (GESIS, 2025a), released in January 2025 and including responses collected between April and September 2023. It features $5,246$ participants and a total of $605$ variables. We work with the ALLBUScompact  [^2] https://search.gesis.org/research_data/ZA8832 (GESIS, 2025b), the openly available version of the ALLBUS, featuring $579$ variables. For data protection reasons, the ALLBUScompact contains information in reduced detail on the socio-demographic characteristics age, origin, occupation and income, and omits all fine-grained geographic variables.
@@
In our preliminary analysis, we explore the effects of using a representative persona collection such as the GGSS Personas—representative of the German population by virtue of mirroring all ALLBUS participants—as well as unrepresentative variants of it, created by systematically oversampling on selected attributes (Section 4).
@@
While it is theoretically feasible to use all $579$ variables featured in the latest release of the ALLBUScompact in the persona prompts, this would unavoidably come at the price of increased context window requirements, computational resources, as well as potentially degrading performance due to the dilution of relevant information or the failure of the model to put equal attention to all variables in a lengthy persona description. For example, Shi et al. (2023) have shown that LLMs can easily be distracted by irrelevant information included in the context window, and de Araujo et al. (2025) have observed performance drops after including irrelevant persona details in persona prompts.
One of the core advantages of general population surveys such as the ALLBUS is the coverage of a variety of different topics, ranging from core demographic variables such as age and gender to very specific, thematic variables covering, e.g., the participant’s political views or lifestyle choices. Therefore, we design a data-driven method that allows us to identify and select the statistically most important variables, instead of having to (artificially) restrict the GGSS Personas to cover only task-specific areas such as socio-demographics, psychological inventories, or political beliefs.
@@
We expect that variables with higher aggregate feature importance score explain more variance of the other variables in the survey. This ranking thus allows us to create versions of the GGSS Personas collection for different computational budgets and needs by selecting only the $k$ most important attributes (TOP-$k$). In situations with restricted prompt length or inference time, versions with less attributes may be preferable, while in unrestricted settings, the number of included attributes may be increased, up to the inclusion of all $380$ survey variables. In our preliminary analysis, we explore the effects of increasing the number of persona attributes on the ability of LLMs to accurately simulate survey response distributions (Section 4).
An important limitation of our approach is its inability to deal with missing values. Because large surveys such as the ALLBUS work with questionnaire splits to manage the number of questions each participant is asked to answer, no participant will have answered all survey questions. Thus, some information is just not available for some of the participants, meaning that for any given variable some persona descriptions will always be incomplete, missing individual attributes. Table 1 shows the extent to which the persona descriptions across the different TOP-$k$ collections are complete. The $k$ thus represents the maximum number of attributes possibly featured in a persona of a given TOP-$k$ collection, and the reported average is the actual number of attributes included. We discuss the implications for the GGSS Personas and our preliminary analysis in Section 6.
@@
There is a broad range of different prompting techniques in the literature Lutz et al. (2025), ranging from the use of key-value pairs in a dictionary or JSON-format (e.g., Castricato et al. (2025)) to the use of full-text persona descriptions (e.g., Ge et al. (2024)), oftentimes generated using LLMs. While personas in the JSON-format are easier to construct and more convenient to work with, the full-text persona descriptions resemble the natural language interactions that especially instruction-tuned LLMs have been optimized for more closely.
We evaluate the impact of using different persona formats for predicting survey response distributions. Since we do not observe a large impact of the persona format on the experimental results, (see Appendix A Figure 7) we use the JSON-format for all following experiments, expecting the results to generalize to other persona formats. However, we make both the JSON as well as the full-text version of the GGSS Personas publicly available. We validate the full-text personas that are LLM-generated from the JSON-personas as input. Two co-authors of this paper independently annotated a sample of $80$ personas ($20$ randomly-sampled personas per TOP-$2$, -44, -88 and -$16$ collection) for hallucinations, misrepresentation or omission of individual attributes. We find that $69$ of $80$ (or 86.2586.25%) full-text personas are accurate depictions of the JSON-personas, with errors occurring rarely through misrepresentation, e.g., of double negations or complex temporal orders, or through the omission of single attributes. We further describe the human validation process in Appendix C.1.